Mathematics 242 – Principles of Analysis Definitions Quiz 1 – September 10, 2004

Your Name:

1. Construct the truth table for ((p ⇒ q) ∧ ∼ q) ⇒ ∼ p)

2. Fill in the blank: A statement that is always true is called a .